OpenStudy (anonymous):

solve the system using a matrix equation: -2x+y+2z=-13 , -x-4y+z=-11 , -y-z=0

OpenStudy (anonymous):

\[\begin{bmatrix} -2&1&2\\ -1&-4&1\\ 0&-1&-1 \end{bmatrix}\begin{bmatrix}x\\y\\z\end{bmatrix}=\begin{bmatrix}-13\\-11\\0\end{bmatrix}\]